OpenStudy (anonymous):

-4√x+9=20 Help Please and Thank you:)!!

OpenStudy (cwrw238):

subtract 9 from both sides -4√x = 11 can you continue from here

OpenStudy (anonymous):

Oh I see what you mean:)

OpenStudy (anonymous):

What would I need to do next?

OpenStudy (anonymous):

-4√x+9=20 -4√x=11 √x=-(11/4) x=√-(11/4) ok ??

OpenStudy (anonymous):

I see! Thank you:)